# AutoCV: Empowering Reasoning with Automated Process Labeling via Confidence Variation

### v2

@@ Appendix C More Related Work
Improving LLMs through learning from feedback has become a prevalent strategy, notably through reinforcement learning from human feedback, which seeks to align LLMs with human values by refining their outputs based on feedback [44, 45, 46]. However, this method faces challenges such as high costs due to manual labor and a lack of real-time feedback capabilities [47, 48]. An alternative strategy involves using self-correcting LLMs, which rely on automated feedback to iteratively adapt and understand the consequences of their actions without heavy reliance on human intervention. This feedback can be derived from inside sources such as the model itself [49, 50] or generation logits [51], and outside sources such as other models [52, 53], tools [54, 55], knowledge bases [56, 57], or evaluation metrics [58, 59].
@@ Appendix D Vanilla Evaluation Methods Description
Given the following question: ’[question]’, and these five answers: 1. [answer] 2. [answer] 3. [answer] 4. [answer] 5. [answer] Which answer is the best? Please analyze and compare the provided answers and then identify the number of the best answers. You should strictly follow the output format requirements and not output any other content. Example: Comparison and Analysis: [analysis]. Best answer: [number]. Let’s Begin!
@@
Appendix E PCH benchmark
In our approach, we employ the LlaMA2-chat model (LlaMA) to generate steps for mathematical reasoning. Through the use of regular expressions, we isolate steps that involve numerical calculations. These steps, specifically the expressions to the left of the "=" sign, are then evaluated using Python’s eval function to verify their correctness against the results on the right-hand side. We denote it as “PCH Detection” and exhibit the results in Table 11.
@@ Appendix E PCH benchmark
Model Pass@5 (%) Self-Consistency (%) PCH Detection (%)
@@
Instances that contain expressions that cannot be evaluated due to unsolvability or incorrect formatting (for example, "x + 1 + 2 = 4") are categorized as non-computational and excluded from the ground truth data. This approach ensures that our ground truth reflects only the model’s computational errors during reasoning, avoiding any overstatement of its computational accuracy.
It’s worth noting that while organizations like DeepMind [15] have undertaken similar annotation tasks with numerous human labelers, referring to them as "trace errors," our method employs Python’s ‘eval‘ function for automated labeling. This strategy provides a scalable and efficient way to approximate trace errors.
To further illustrate how we calculate $\Delta_{conf}^{t}$ to identify PCH error, consider the following example. As shown in Figure 2, the highlighted errors in red signify hallucinations during the reasoning process. This is detected by the model’s confidence decreasing significantly.
@@ F.1 Training Hyperparameters
Our experiments were conducted in a computing environment equipped with 8 NVIDIA A100 GPUs, each having 40GB of memory. All models are fine-tuned in a full-parameter setting.
We employed the AdamW optimizer for model training over 1 epochs, with a batch size of 512. The learning rate was set at $2e\times 10^{-6}$, incorporating a 3% learning rate warmup period. Below, we present a comprehensive overview of the training hyperparameters utilized. These parameters were consistently applied across training both process-supervised and outcome-supervised methods in our experiments in Table 12.
@@
Before we train a verifier, it is essential to first train a supervised fine-tuning model (SFT). This SFT model enables us to generate responses and label outcome supervision data by checking whether the final answer being correct or not. The training parameters for the SFT model are outlined in Table 13. For further training details regarding OSV, refer to [22].
@@ F.2 Generation Settings
For the generation of results using the "greedy" strategy, we set the temperature parameter to 0.0 and 0.7 for the "pass@k" strategy. To present unbiased results for "pass@k", we follow the calculation method outlined in [43]. Specifically, we generate $n=20$ samples for each instance, evaluate the number of correct samples passing unit tests, and then calculate the unbiased estimator for pass@k. We repeat the above experiments 5 times to obtain the 95% confidence intervals reported in the Section 5.1
@@ F.3 Process-Supervision Labelling Strategy
We detail the implementation process for both our proposed method (Process Ours) and the comparative method (Process MCS). Initially, we train a generator through standard supervised fine-tuning using training datasets from GSM8K and MATH, comprising approximately 15,000 instances.
Following this, we utilize the trained generator to produce 10 samples for each unlabeled prompt within the training datasets. These samples are then labeled as either 0 or 1 based on the accuracy of the final answer obtained.
For Process Ours, we subsequently train an output-supervised verifier using the total of 15,000 * 10 generated samples. These samples are then relabeled using a relative confidence change criterion (refer to Eq. 3), followed by retraining a Process Ours-supervised verifier.
In the case of Process MCS, each of the 15,000 * 10 samples is divided into reasoning paths. For each incomplete reasoning path, we employ the generator to generate eight complete reasoning paths. The accuracy of the final answer is then computed for each complete path, and the fraction of correct paths is utilized to relabel the samples. Finally, a Process MCS-supervised verifier is retrained.
Note: We do not sample 50 times for each training problem as we do in the main results, since 50 * 15,000 data points for MCS labeling would take approximately 10-12 days.

### v4

@@ Appendix C More Related Work
Improving LLMs through learning from feedback has become a prevalent strategy, notably through reinforcement learning from human feedback, which seeks to align LLMs with human values by refining their outputs based on feedback [43, 44, 45]. However, this method faces challenges such as high costs due to manual labor and a lack of real-time feedback capabilities [46, 47]. An alternative strategy involves using self-correcting LLMs, which rely on automated feedback to iteratively adapt and understand the consequences of their actions without heavy reliance on human intervention. This feedback can be derived from outside sources such as other models [48, 49], tools [50, 51], knowledge bases [52, 53], evaluation metrics [54, 55] or generation logits [56].
@@ Appendix D Vanilla Evaluation Methods Description
Given the following question: ’[question]’, and these five answers: 1. [answer] 2. [answer] 3. [answer] 4. [answer] 5. [answer] Which answer is the best? Please analyze and compare the provided answers and then identify the number of the best answer. You should strictly follow the output format requirements and not output any other content. Example: Comparison and Analysis: [analysis]. Best answer: [number]. Let’s Begin!
@@
Appendix E Math Calculation Error Benchmark
Methodology
We utilize the LlaMA2-chat model (LlaMA) for mathematical reasoning step generation. Using regular expressions, we extract computational steps and evaluate expressions to the left of the "=" sign using Python’s eval function to verify their correctness against right-hand side results. We term this process "Math Calculation Error Detection" and present the results in Table 12.
@@ Appendix E Math Calculation Error Benchmark
Model Pass@5 (%) Self-Consistency (%) Math Calculation Error Detection (%)
@@
Data Processing
Non-computational expressions (e.g., "x + 1 + 2 = 4") that cannot be evaluated due to unsolvability or incorrect formatting are excluded from the ground truth data. This ensures our ground truth accurately reflects only computational errors during reasoning. While DeepMind [15] employs human labelers for "trace errors" annotation, our automated approach using Python’s eval function provides a scalable alternative.
Error Detection Example
To demonstrate our calculation of $\Delta_{conf}^{t}$ for math calculation error identification, we present an example in Figure 2. The red highlights indicate calculation errors in the reasoning process, detected through significant decreases in model confidence.
@@ F.1 Training Hyperparameters
Computing Infrastructure
Our experiments were conducted using 8 NVIDIA A100 GPUs, each with 40GB of memory. All models underwent full-parameter fine-tuning using the AdamW optimizer.
Verifier Training Configuration
For both process-supervised and outcome-supervised methods, we maintain consistent training parameters as detailed in Table 13. The training process spans 1 epoch with a batch size of 512 and a learning rate of $2\times 10^{-6}$, incorporating a 3% learning rate warmup period.
@@
SFT Model Prerequisites
Prior to verifier training, we establish a supervised fine-tuning (SFT) model. This model serves to generate responses and facilitate outcome supervision labeling by validating final answer correctness. The SFT model’s training parameters are specified in Table 14. For comprehensive details regarding OSV implementation, refer to [22].
@@ F.2 Generation Settings
Parameter Settings
We employ two distinct temperature configurations: temperature = 0.0 for the greedy decoding strategy and temperature = 0.7 for the pass@k evaluation strategy.
Pass@k Evaluation Protocol
To ensure unbiased estimation of pass@k metrics, we follow the methodology established in [41]. We generate $n=20$ samples per problem instance and evaluate each sample against unit tests to determine correctness. The unbiased estimator for pass@k is then computed based on these results. This entire process is repeated 5 times to establish 95% confidence intervals, as detailed in Section 5.1.
Score Aggregation Methodology
Our primary aggregation mechanism employs a multiplicative approach for step-level scores. This method computes the product of confidence scores across all steps, yielding a compound probability that represents the overall likelihood of solution correctness.
Comparative Analysis of Aggregation Functions
Following [16], we conducted a comparative analysis of different aggregation strategies. While our primary approach utilizes the product of step-level scores, we also investigate the minimum of step-level scores as an alternative approach. The comparative results for these aggregation functions on the GSM8K dataset are presented below:
Table 15: Comparison of aggregation functions on GSM8K dataset.
Response Generator Product Minimum
Mistral-Instruct 60.72 61.17
Mixtral-Instruct 74.07 72.45
Qwen 85.00 84.28
As shown in Table 15, the choice of aggregation function can have a slight impact on the results, with the product function generally performing comparably or slightly better than the minimum function across different response generators.
@@ F.3 Process-Supervision Labelling Strategy
Overview
Our implementation encompasses two distinct labelling methodologies: our proposed approach (Process Ours) and a comparative Monte Carlo Tree Search approach (Process MCTS).
Implementation Procedure
The procedure consists of the following sequential stages:
  1. 1.
    Initial Generator Training: We conduct supervised fine-tuning of a generator using a combined dataset of approximately 15,000 problems from GSM8K and MATH.
  2. 2.
    Sample Generation: The trained generator produces 10 distinct solution attempts for each unlabeled prompt within the training datasets. Each solution receives a binary label (0 or 1) based on the correctness of its final answer.
  3. 3.
    Method-Specific Processing:
    1. (a)
      Process Ours:
      • Train an output-supervised verifier using the complete set of 150,000 generated samples (15,000 × 10)
      • Relabel samples using a relative confidence change criterion (as defined in Equation 3)
      • Retrain the verifier using these refined labels
    2. (b)
      Process MCTS:
      • Decompose each of the 150,000 samples into constituent reasoning paths
      • For each incomplete reasoning path, generate eight complete path variations
      • Compute the accuracy for each complete path
      • Calculate the proportion of correct paths to determine the final label
      • Retrain the verifier using these MCTS-derived labels
Technical Note
While our main experimental results employ 50 samples per problem, we restrict this implementation to 10 samples due to computational constraints. The full MCTS labeling process for 50 × 15,000 samples would require approximately 10-12 days of computation time.